Scores of an achievement test show that it follows a normal distribution. Its mean is 79 with a



standard deviation of 8. Find the probability that students’ scores will be higher than the passing score



of 70.

1
Expert's answer
2022-06-14T00:11:27-0400

We have a normal distribution, "\mu=79, \sigma=8."

Let's convert it to the standard normal distribution,

"z=\cfrac{x-\mu}{\sigma}=\cfrac{70-79}{8}=-1.13;"

"P(X>70)=P(Z>-1.13)=1-P(Z<-1.13)=\\\\\n=1-0.1292=0.8708\\ \\text{\u200b(from z-table).}"